A 4.0-kg block travels around a 0.50-m radius circle with an angular velocity of 12 rad/s. Its angular momentum about the center of the circle is:


a: 12 kg.m2 /s


b. 24 kg.m2 /s


c. 48 kg.m2 /s


d. 6 kg.m2/s


1
Expert's answer
2021-12-20T10:28:04-0500

Given:

"m=4.0\:\rm kg"

"r=0.50\:\rm m"

"\omega=12\:\rm rad/s"


The angular momentum is given by

"L=I\omega=mr^2\omega"

"L=4.0*0.50^2*12=12\:\rm kg\cdot m^2/s"